Solve the problem with complete solutions. Determine the amount of heat released by a 3.0 kg of water when it cools from 80 degree Celsius to 10 degree Celsius.


1
Expert's answer
2021-12-29T12:35:21-0500

The amount of heat released is


"Q=cm(t_f-t_i),\\
Q=4200·3·(10-80)=-882000\text{ J}."